Two identical positive point charges (q1 = q2 = +2.00 nC) are placed at the bottom two vertices of an equilateral triangle of edge length 5.00 cm. One negative point charge (q3= -3.00 nC) is placed at the apex. (a) Calculate the electrical potential produced by the combination of all three point charges together at the midpoint of the base of the triangle. (Remember that each angle in an equilateral triangle is 60 degrees.) (Answer: 816 V) (b) Calculate the amount of work that would need to be done in order to bring a point charge q4 = +1.00 nC in from infinity and place it at the midpoint of the base.
Answer #1

(a)

the formula for potential due to point charge is

V = kq/r

V total = kq1/ r1 + k q2/ r2 + kq/ r3

r3 = sqrt ( (0.05)^2 - ( 0.025)^2

=0.0433 m

V total = 9 * 10^9 * 10^-9 ( 2/0.025 +2/0.025 - 3/0.0433)

=816 V

(b)

W = q V total = 1* 10^-9 (816) = 8.16* 10^-7 J